The brakes of a 1000 π‘˜π‘” automobile can exert a retarding force of 3,000𝑁. Find the distance the car will move before stropping if it is traveling at the rate of 24 π‘š/𝑠 when the brakes are applied (96 π‘š).Β 


1
Expert's answer
2021-10-03T13:22:05-0400

"a=F/m=3000/1000=3\ (m/s^2)"


"s=v^2/(2a)=24^2/(2\cdot3)=96\ (m)" . Answer